OpenStudy (anonymous):

find derivative of f(x)=(lnt +t)^(1/2)

OpenStudy (anonymous):

you have this f(x)=(lnt +t)^(1/2) dy/dx = 1/2((lnt +t))^(-1/2) . d/dx((lnt +t)) do it now

OpenStudy (anonymous):

I got that far on my own, my problem is simplifying the next step ... f'(x)=1/2(in t+t)^(-1/2) *(1/t+1)

OpenStudy (walters):

did u get something like

OpenStudy (anonymous):